lss6 tt60 beam commissioning on 24 th august 2006
Download
Skip this Video
Download Presentation
LSS6/TT60 beam commissioning on 24 th August 2006

Loading in 2 Seconds...

play fullscreen
1 / 14

LSS6/TT60 beam commissioning on 24 th August 2006 - PowerPoint PPT Presentation


  • 56 Views
  • Uploaded on

LSS6/TT60 beam commissioning on 24 th August 2006. On behalf of E.Carlier, E.Gaxiola, M.Gourber-Pace, E.Hatziangeli, L.Jensen, R.Jones, V.Kain, A.Koschik, M.Lamont, J.Uythoven, J.Wenninger and many more….

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' LSS6/TT60 beam commissioning on 24 th August 2006' - yosefu


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
lss6 tt60 beam commissioning on 24 th august 2006

LSS6/TT60 beam commissioning on 24th August 2006

On behalf of E.Carlier, E.Gaxiola, M.Gourber-Pace, E.Hatziangeli, L.Jensen, R.Jones, V.Kain, A.Koschik, M.Lamont, J.Uythoven, J.Wenninger and many more…

http://proj-lti.web.cern.ch/proj-lti/LTIcoordination/relatedmeetings/Extraction6/Extraction-test.htm

LTC 13th September 2006

overview
Overview
  • New fast extraction channel in LSS6 commissioned with beam
    • LHC beam 1 through TT60 towards TI2: ~100 m of TT60 onto TED beam dump.
  • Successful and efficient test
    • Beam extracted to TED on first attempt
    • Equipment all worked, eventually, almost perfectly.
    • Controls & SW all worked, almost perfectly.
    • Almost all planned measurements completed in ~16 hours.
    • ~3500 pilot extractions with 5 109 p+, 12 extractions with 12 bunches, 5 1011 p+.
    • No major issues discovered.

First shot on BTV.610317 (TED), 12:35:03

LTC 13th September 2006

test program overview
Test program & overview
  • PREPARATION………………………………..
  • CYCLE & EQUIPMENT SETTING UP………
  • LHC BEAM SETTING UP IN SPS…………..
  • PILOT EXTRACTION………………………...
  • APERTURE CHECKS………………………..
  • TPSG ALIGNMENT CHECKS……………….
  • STEERING AND OPTICS CHECKS………..
  • 3 of 4 KICKER TANK TEST………………….
  • KICKER WAVEFORM MEASUREMENT…..
  • REPRODUCIBILITY………………………….
  • MULTI-BUNCH EXTRACTION………………
  • RECOVERY……………………………………

OK

OK

OK

OK

OK

OK

Partial (during 5 and 6)

OK

OK

Skipped

OK

OK

LTC 13th September 2006

how the time was used
How the time was used

Intermediate extraction ~0.2 h

PilotIntermediate ~1 h

Recovery ~1.5 h

Preparation ~4.5 h

Pilot extraction ~9 h

  • Total ~1.751013 p+ on TED (cf 51013 maximum announced)
  • About 3500 extractions made using fast LHC cycle

LTC 13th September 2006

extraction and aperture measurements
Extraction and aperture measurements
  • Verified positions of all extraction elements with circulating or extracted beam
    • Detailed data reduction still in progress but no major surprises
    • Same measurements will be made for LHC extraction septa MSD and TCDS
  • Calibrated beam loss measurements and profile for circulating and extracted beams
    • Different signatures along the extraction channel to LSS4, as expected from TPSG designs
    • Needed to set BLM interlock thresholds
    • Allows subsequent quantification of absolute loss levels
  • Measured aperture along TT60 with varying corrector/septum strengths
    • Handicapped by lack of BPCKs and fact that BCT is halfway between extraction and TED
    • Used positions on screens to estimate maximum possible excursions
    • Results so far analysed fit well with expected aperture from model (madx plus physical layout)

V.Kain

LTC 13th September 2006

optics checks
Optics checks
  • Using data taken for aperture and extraction calibration measurements
  • Only BTV data available at that time (BPCK firmware problem was being solved)
  • Automatic data analysis difficult
    • Positions not returned consistently by BTV device
    • SDDS used to get positions from 2-D profiles
  • Already see some interesting results: reasonably good agreement with MAD model
  • Will extend to all corrector/kicker/septum excitations and measured responses
    • Preparing input files from raw data for thorough analysis – rather painstaking

V.Kain/A.Koschik

LTC 13th September 2006

mke results saturation calibration
MKE results – saturation/calibration
  • Varied MKE strength from nominal (26 kV) to 33% higher (35 kV).
  • Checked for saturation effects at higher current /field (up to 0.13 T).
  • No evidence of non-linearity seen over the range tested
  • Would allow use of 3 MKE for extraction – merits being discussed

J.Uythoven/E.Gaxiola

LTC 13th September 2006

mke results waveform
MKE results - waveform
  • Varied MKE delay wrt beam position to measure kicker waveform
    • as planned for MKD/MKB systems during commissioning of LHC dump system

Bunches being kicked on the rising edge, making 1 turn and being extracted

  • Find ±1% ripple
    • ±0.5% specification – needs improving
  • Flat-top about 18 ms
    • Too long – needs shortening
    • Bunches kicked outwards on rising edge and make full SPS turn before extraction
  • See effect of non-linear QFA618 field
    • Should measure this explicitly

±1%

J.Uythoven/E.Gaxiola

LTC 13th September 2006

beam instrumentation specifics
Beam Instrumentation specifics
  • Useful for AB/BI in verifying new electronics in place for LHC beam extraction
  • Found with beam:
    • Some cabling/SW inversions in LSS60 BPCEs were found with the beam
    • Firmware problem with BPCKs in TT60
  • Difficulties in accurate measurement of low intensity with TT60 BCT (kicker noise)
  • Certain problems remain to be solved:
    • Most difficult probably signal to noise ratio for the very low intensity pilot beam
    • Problem with BPMI firmware been investigated and fix believed available.
    • possible BTV vs BPCK scaling errors being investigated
  • BTVs had x10 smaller RMS (31 mm) in position measurement than BPCKs!

L.Jensen

LTC 13th September 2006

controls sw specifics
Controls & SW specifics
  • Major progress made since earlier tests
  • All CO applications ran smoothly
    • LSA AP (EquipState, Cycle Manager )
    • OASIS Viewer
    • Fixed Displays
    • DB Logging
    • SDDS Logging
    • ALARMS
  • Some adjustments performed on the fly to Fixed Displays and DB logging
  • Main outstanding issues concern logging
    • Systematic wrong data time stamping at source for SPS Ring BLM + MUGEF
    • Septa logged asynchronously (NO timing in FE)
    • Sporadic holes in data
    • OASIS scope and channel settings
  • Missing alarms from mugefs

Fixed display showing all 7 LSS6/TT60 BTVs

M.Gourber-Pace

LTC 13th September 2006

radiation protection
Radiation protection
  • Continuous monitoring of dose rate at accessible gates showed background levels during test – max. predicted was 1 mSv/h (pessimistic)
  • Radiation survey showed activation levels as expected – 450 mSv/h at TED
  • Waiting for results from dosimeters

Heinz Vincke/N.Conan

LTC 13th September 2006

positive aspects
Positive aspects
  • High beam availability from SPS
  • Fast 8.4 s LHC cycle and pilot bunch from SPS  rapid data taking
  • Controls & SW very stable and mature
    • Reaping the LSA harvest
  • “Fast” switch to 12 bunches at midnight (~1 hour)
  • Very few problems with equipment systems
    • Most problems shaken out in dry runs in preceding weeks
  • Comprehensive and stable logging
  • Excellent support from OP, BI and CO

LTC 13th September 2006

problems issues encountered
Problems/issues encountered
  • No time to condition MKE kickers to full voltage prior to the test
    • Sparks occurred in magnets - caused several vacuum interlocks
  • SPS BPCEs with polarity/plane inversions
    • Could have been seen before the test, with another beam
  • TT60 BPCKs needed several hours’ from experts to work correctly
    • Not available for aperture/response measurements
  • Not possible to change polarity of MDLV for aperture tests
    • Still some LSA issues with polarity change and ‘zero’ current functions
  • Synchronisation of logged data imperfect
    • TIMESTAMP data different within cycle from different equipment (BPCK, BTV, mugef)
    • Data 2 cycles late for BLMS!
    • Makes data analysis much more difficult: has to be properly addressed for 2007!!!

LTC 13th September 2006

outlook for 2006 2007
Outlook for 2006/2007
  • Some measurements still to make in interleaved LSS4/6 test
    • Shorten MKE pulse and use only 3 kickers (tbc), and measure new waveform
    • Measure of non-linear field in QFA618
    • Full optics/response measurement with BPCKs
    • Repeat subset of aperture measurements, with BPCKs available
  • Commissioning still to make in 2007
    • Interlock system available and to be tested
    • High intensity tests - full LHC beam extracted
    • Interleaved low and high intensity extraction with LSS4 and LHC injection sequencing
    • Get ready for TI2 commissioning later in the year

LTC 13th September 2006

ad